# RacerF: Lightweight Static Data Race Detection for C Code

### v1

@@ 8 Implementation and Experimental Evaluation
We have implemented our approach as a plugin of the Frama-C platform called RacerF  [^4] RacerF is available under the MIT license at https://github.com/TDacik/Deadlock_Racer as a part of our bigger project for detection of concurrency bugs.. The individual analyses are written as OCaml functors parametrised by a module implementing the backend analysis. We also provide a Python wrapper which runs a combination of under- and over-approximating analyses and also performs preprocessing (currently, the preprocessing is implemented outside of our analyser for technical reasons, thus the analyser is slowed a bit by the need of repeated parsing). To solve the systems of equations from Sections 5.2 and 5.3 we use the chaotic iteration strategy with widenings [6] implemented in the Ocamlgraph library [8]. The widening operators are those provided by the $\mathsf{CValue}$ domain.
@@ 8.1 SV-COMP Benchmarks
The SV-COMP sub-category $\mathsf{NoDataRace\!-\!Main}$ is, to our knowledge, the largest publicly available data race benchmark. Many of its programs are small but intricate tests provided by the participants. Our main goal is to demonstrate the scalability of our approach and to also show that its heuristic nature does not compromise its precision even on those intricate examples. The benchmark also contains implementations of lock-free algorithms, which are out of the scope of our tool. We have implemented a single heuristic that detects active waiting (essentially a loop with an empty body) as a characteristic pattern for lock-free algorithms, and when it is present, returns an unknown result.
We compare with the five best-performing tools participating in SV-COMP 2025 [14, 10, 18, 26] (including RacerFsv-comp, an older version of RacerF) and we also include RacerFunder and RacerFover representing RacerF running in under- and over-approximating mode only, respectively. We wanted to include also the O2 checker [21], but since its results were not good and all the other tools were configured specifically for this benchmark by their authors, we do not consider such a comparison fair.
Table 1: Results for the SV-COMP sub-category NoDataRace-Main. The columns represent true negatives, false positives, true positives, false negatives, out-of-resource (time or memory), and times on local and competition machine, respectively.
@@
Analyser  TN  FP  TP     FN     RO    Score  Time1*    Time2**
RacerF 674 4 110 0 0 1394 4480 -
RacerFover 674 6 87 0 0 1339 3810 -
RacerFunder 680 4 110 6 0 1208 2630 -
@@
Goblint 712 0 0 0 3 1424 - 5200
Deagle 685 0 203 1 43 1541 - 44k
GemCutter 578 0 161 0 263 1317 - 290k
Automizer 609 0 121 0 274 1339 - 170k
The results are summarised in Table 1. Unlike the competition results [1], we present the results ignoring the witness validation. RacerF produced 4 false alarms, all of them on programs where a data race is ruled out by some intricate condition. Two of them are variants of a program where a data race does not manifest because of an array being searched in parallel contains only zeros and a certain condition is thus triggered just once (which is the kind of behaviour that our tool cannot handle).
@@
Interestingly, running RacerF in either the over- or under-approximating mode does not compromise its precision too much, leading to an additional 2 false positives (and 23 must-races classified just as may- and thus not reported) and 6 false negatives, respectively. However, we are aware that there are more programs where approximation could make a difference, but RacerF returns an unknown result for another reason.
Compared to participants of SV-COMP’25, RacerF placed second in the competition ranking and third in our evaluation. Compared to Goblint, which never reports races, RacerF can produce correct results for more programs. Compared to Deagle, which benefits from the fact that many programs are small and their loops can be fully unrolled, no version of RacerF runs out of time or memory.
Where RacerF shines is the $\mathsf{ldv\!\!-\!\!linux\!\!-\!\!3.14\!\!-\!\!races}$ benchmark subset containing 5 programs from the Linux kernel (unfortunately, all of them are race-free). None of the participants reaching a positive score can provide correct verdicts for those programs (except Goblint which handles one of them), but RacerF can correctly analyse each of them under two minutes.
Table 2: Results for student programs. The columns are true positives, false negatives, no race reported, race reported, out-of-resources and time in second.
@@
Analyser  TP  FN  NR     RR    OR    Time
RacerF 16 2 47 0 0 406
O2 7 21 87 0 0 137
Goblint 27 2 47 40 0 39.4
Deagle 2 1 5 0 80 6300
@@ 8.2 Student Programs
To test our tool on more non-crafted programs, we performed an evaluation on a set of 116 student programs from [16]. The programs are rather short (around 100 – 300 LOC) and implement the same heavily concurrent algorithm parametric in the number of threads, but they differ in various intricate bugs (for their detailed discussion see [16]). We have identified 28 data races using dynamic analysers. However, there may be more that those tools missed. One additional race was discovered by all RacerF, O2, and Goblint. The issue could manifest only in a very rare situation when two threads simultaneously failed and called a clean-up function that assigns a global pointer without synchronization. The time limit was set to 60 seconds and the memory limit to 1GB.
We have compared with Goblint (in version from SV-COMP’25 but with the default configuration as the other available configurations did not improve the results and led to timeouts) and Deagle (version from SV-COMP’25) that outperformed us in SV-COMP and with O2 (implemented in Coderrect [^5] The link https://coderrect.com from which we got ver. 1.1.3 is no longer available.) that we selected as a recent representative of lightweight and scalable race detection tools.
The results are given in Table 2. Deagle ran out of the resources on most of the programs, providing just a few results. Interestingly, Goblint and RacerF claimed the same 47 programs as race-free. The two missed races by RacerF and Goblint are also the same and both are related to problems with lock initialisation that the tools do not support. A detailed inspection suggests that both tools struggle with programs that use an array where each thread accesses only the element at its dedicated index.
@@ 9 Conclusions and Future Work
We have presented RacerF and showed that it achieves a nice compromise between precision and scalability both on crafted benchmarks and real-world programs. In the future, we would like to target other classes of concurrency bugs. We already have an implementation for deadlock detection based on top of our lockset analysis, which is not discussed in this paper. Another interesting direction is combination with dynamic approaches, e.g., guiding insertion of noise injection by results of static analysis.

### v2

@@ 8 Implementation and Experimental Evaluation
The RacerF  [^4] RacerF is available under the MIT license at https://github.com/TDacik/Deadlock_and_Racer. analyser based on the above presented principles is implemented as a plugin of the Frama-C platform as a part of our bigger project for detection of concurrency bugs. The individual analyses are written as OCaml functors parametrised by a module implementing the backend analysis. We also provide a Python wrapper which runs a combination of under- and over-approximating analyses and also performs preprocessing (currently, the preprocessing is implemented outside of our analyser for technical reasons, thus the analyser is slowed a bit by the need of repeated parsing). To solve the systems of equations from Sections 5.2 and 5.3 we use the chaotic iteration strategy with widenings [7] implemented in the Ocamlgraph library [9]. The widening operators are those provided by the $\mathsf{CValue}$ domain.
@@
We have conducted a series of experiments to support the following two hypotheses:
  • Despite having no formal guarantees, RacerF can efficiently analyse intricate programs without sacrificing a precision. We validated this hypothesis on a collection of benchmarks from SV-COMP containing 1029 programs, many of them being intricate examples and corner cases contributed by participants to showcase capabilities of their tools.
  • RacerF can detect data races in real-world programs for which other existing static analysers do not provide precise results (either because of too coarse approximation or running out of resources). We validated this hypothesis on a set of student programs from [17].
@@ 8.1 SV-COMP Benchmarks
The SV-COMP sub-category NoDataRace-Main is, to our knowledge, the largest publicly available data race benchmark. Many of its programs are small but intricate tests provided by the participants. Our main goal is to demonstrate the scalability of our approach and to also show that its heuristic nature does not compromise its precision even on those intricate examples. The benchmark also contains implementations of lock-free algorithms, which are out of the scope of our tool. We have implemented a single heuristic that detects active waiting (essentially a loop with an empty body) as a characteristic pattern for lock-free algorithms, and when it is present, returns an unknown result.
We compare with the five best-performing tools participating in SV-COMP 2025 [20, 21, 22, 28] (including RacerFsv-comp, an older version of RacerF) and we also include RacerFunder and RacerFover representing RacerF running with under- and over-approximating thread analysis only, respectively. We wanted to include also the O2 checker [25], but since its results were not good and all the other tools were configured specifically for this benchmark by their authors, we do not consider such a comparison fair.
Table 1: Results for the SV-COMP sub-category NoDataRace-Main. The columns represent true negatives, false positives, true positives, false negatives, out-of-resource (time or memory), score (as defined by SV-COMP rules, but excluding witness validation), and times (excluding timeouts) on local and competition machine, respectively.
@@
Analyser TN FP TP FN RO Score Time1 [s] Time2 [s]
[…]
RacerFover 674 6 83 0 0 1335 3970 -
[…]
@@
[…]
Deagle 685 0 203 1 43 1541 - 10k
[…]
UAutomizer 609 0 121 0 274 1339 - 120k
The results are summarised in Table 1. Unlike the competition results [3], we present the results ignoring the witness validation. RacerF produced 4 false alarms, all of them on programs where a data race is ruled out by some intricate condition. Two of them are variants of a program where a data race does not manifest because an array being searched in parallel contains only zeros and a certain condition is thus never triggered which makes the potential racy access unreachable (which is the kind of behaviour that our tool cannot handle). The other two contain locking patterns that are beyond the scope of our lockset analysis.
@@
Interestingly, running RacerF in either the over- or under-approximating mode does not compromise its precision too much, leading to an additional 2 false positives (and 22 must-races classified just as may- and thus not reported) and 6 false negatives, respectively.
Compared to participants of SV-COMP’25, RacerF placed second in the competition ranking and third in our evaluation. Compared to Goblint, which never reports races, RacerF can produce correct results for more programs. Compared to model checker Deagle, which benefits from the fact that many programs are small and their loops can be fully unrolled, no version of RacerF ever runs out of time or memory.
Where RacerF shines is the $\mathsf{ldv\!\!-\!\!linux\!\!-\!\!3.14\!\!-\!\!races}$ benchmark subset containing 5 programs derived from the Linux kernel (all of them are race-free). None of the participants with a positive score can provide correct verdicts for those programs (except Goblint which handles one of them), but RacerF can correctly analyse each of them under two minutes.
Table 2: Results for student programs. The columns are true positives, false negatives, no race reported, race reported, out-of-resources, and time (excluding timeouts).
@@
Analyser  TP  FN  NR     RR    OOR    Time [s]
[…]
O2 7 23 86 0 0 54
[…]
Goblint 27 3 46 40 0 41
[…]
@@ 8.2 Student Programs
To test our tool on more non-crafted programs, we performed an evaluation on a set of 116 student programs from [17]. The programs are rather short (around 100 – 300 LoC) and implement the same heavily concurrent algorithm parametric in the number of threads, but they differ in various intricate bugs (for their detailed discussion see [17]). We have identified 29 data races using dynamic analysers. However, there may be more that those tools missed. One additional race was discovered by all RacerF, O2, and Goblint. The issue could manifest only in a very rare situation when two threads simultaneously failed and called a clean-up function that assigns a global pointer without synchronization. The time limit was set to 60 seconds and the memory limit to 1GB.
We have compared with Goblint  [^5] We use Goblint in the version from SV-COMP’25 but with the default configuration as the other available configurations did not improve the results and led to timeouts. Further, to provide more detailed results, we run Goblint via its SV-COMP wrapper that never reports detected races (Goblint ${}_{\!\mathrm{sv}}$) as well as without it. and Deagle (version from SV-COMP’25) that performed better than us in SV-COMP, and with O2 (implemented in Coderrect [^6] The link https://coderrect.com from which we got version 1.1.3 is no longer available.) that we selected as a recent representative of lightweight and scalable race detection tools.
The results are given in Table 2. As expected, since all the programs contain unbounded loops and also an unbounded number of threads is created, Deagle ran out of the resources on almost all programs and provided an answer in only 10 cases. Interestingly, Goblint and RacerF claimed the same 46 programs as race-free (but RacerF without providing any false positives). The three missed races by RacerF and Goblint are also the same and are caused by an undefined behaviour related to re-initialisation and destroying of mutexes that still may be used by some threads (neither of the tools supports detection of such problems). A detailed inspection suggests that both tools struggle with programs that use an array where each thread accesses only the element at its dedicated index.
@@ 9 Conclusions and Future Work
We have presented a novel data race detector RacerF and performed a series of experiments indicating that RacerF achieves a nice compromise between precision and scalability both on crafted benchmarks and real-world programs. In the future, we would like to target other classes of concurrency bugs. We already have an implementation for deadlock detection based on top of our lockset analysis, which is not discussed in this paper. Another interesting direction is combination with dynamic approaches, e.g., guiding insertion of noise injection by results of static analysis.